OpenStudy (anonymous):

Which pair of angles has congruent values for the sin x° and the cos y°? (5 points) Select one: a. 35°; 55° b. 35°; 145 c. 35°; 70° d. 35°; 35°

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@campbell_st @robtobey @RadEn

OpenStudy (raden):

trials one by one. a. 35° ; 55° check it, is the value of sin 35° = cos 55° ??

OpenStudy (anonymous):

yes :)

OpenStudy (raden):

you can use the identity : sin (a) = cos(90 -a) or cos (a) = sin (90 -a)

OpenStudy (anonymous):

thnxs:)

OpenStudy (raden):

sin(35)° = cos(90 - 35) = cos 55° directly, it is be right :)